OpenStudy (haleyelizabeth2017):

What is the equation for the inverse of y=1/2 sin x?

OpenStudy (phi):

swap x and y , and then solve for y \[ x= \frac{1}{2} \sin y\]

OpenStudy (haleyelizabeth2017):

Oh really? That's it?

OpenStudy (lolacole12):

Technically, y = (1/2)sin(x) does NOT have an inverse function because y = (1/2)sin(x) fails the horizontal line test; also, this should be apparent because sin(x) is periodic. y = arcsin(x) is not the true inverse of y = sin(x); rather, y = arcsin(x) the inverse of y = sin(x) on the interval [-π/2, π/2], the interval that contains x = 0 and is strictly increasing. Now, if you want to find the inverse for y = (1/2)sin(x) when x is on [-π/2, π/2], then just switch x and y and re-solve for y. Switching x and y gives: x = (1/2)sin(y) ==> sin(y) = 2x. Taking the arcsine of both sides gives: y = f^-1(x) = arcsin(2x).

OpenStudy (haleyelizabeth2017):

So it's basically the same thing I did with something else just like last month x'D Awesomeness

OpenStudy (lolacole12):

lol

OpenStudy (haleyelizabeth2017):

Never realized it was that easy lmao

OpenStudy (lolacole12):

so you good or another question

OpenStudy (haleyelizabeth2017):

Some of this trig stuff I think I just overthink

OpenStudy (haleyelizabeth2017):

I'm good for now, if I have another question, I'll open a new post

OpenStudy (lolacole12):

kk

OpenStudy (haleyelizabeth2017):

ty :)